# Grid-Based Continuous Normal Representation for Anomaly Detection

### v1

@@ 0.B.2 Generalization Ability
We have shown the effectiveness of the continuous feature space when sufficient training data is available. However, in this subsection, we suppose a more challenging circumstance where the quantity of normal training data is significantly limited. When the normal images are not sufficient, models often suffer from learning the representative characteristics of the class. For instance, in discrete feature space, if an input feature significantly deviates from the normal features in the space, the model struggles to distinguish whether the distance stems from unseen normal or abnormal features since it has few normal features to be referenced. As shown in Tab. 7, PatchCore, which adopts discrete feature space, shows limited performance under these constraints. On the other hand, GRAD allows continuous normal representation by interpolating the learned normal features in spite of the scarcity, resulting in superior performance under few-shot settings. This demonstrates the generalization ability of the continuous space even in a more challenging scenario. Despite outperforming the other baselines, GRAD’s performance under the extremely scarce dataset (i.e., 1-shot scenario) leaves room for improvement and we believe is not entirely satisfactory for practical uses. We recognize this as a limitation of GRAD, which will be discussed in Appendix 0.C.
@@
PatchCore UniAD GRAD PatchCore UniAD GRAD
1-shot 82.7 80.3 82.3 91.8 91.2 92.2
2-shot 87.5 82.2 89.0 93.8 92.4 95.4
4-shot 89.3 83.6 92.1 94.6 93.1 96.3
@@
Table 8: Ablation study on the grid dimension ($C_{l}$ and $C_{g}$) of local and global representations. The values in the parenthesis denote the grid resolution ($R_{l}$ and $R_{g}$) of each dimension. For instance, 2D (4) in the first row denotes that we set the grid dimension to 2 and each dimension’s resolution to 4. The performances are measured in AUROC on MVTec AD. GRAD’s setting is highlighted in bold.
@@ 0.B.3 More Ablation Studies
Grid dimension. We implemented GRAD with different grid dimensions (equivalent to the dimension of coordinates). As shown in Tab. 8, although the higher dimension of local and global representation slightly increases the detection and localization performance respectively, two representations both with 2-dimension are well balanced, resulting in the optimal performance. However, the results indicate GRAD is not sensitive to grid dimensions, which can be beneficial in a practical scenario.
Feature refinement. We conducted an ablation study on the feature refinement of GRAD, which employs both Mean Squared Error (MSE) and cosine similarity. As shown in Tab. 9, MSE and cosine similarity show comparable performance when they are applied individually. By combining the two metrics, GRAD achieves the best performance.
@@
0.B.4 More Quantitative Results
Fig. 6 illustrates more qualitative results of GRAD.
Figure 6: Qualitative results of GRAD for each class of MVTec AD. Each row represents anomaly images, ground truths, prediction of UniAD, and prediction of GRAD.
@@ Appendix 0.C Limitation
Although GRAD demonstrates its superior generalization capabilities compared to existing methods, as shown in Sec. 0.B.2, it struggles when dealing with novel features. For instance, GRAD can generalize a sparse set of features in a class (i.e., few-shot) by interpolating within their feature space. However, this cannot be the case with extremely limited or no data (i.e., 1-shot or zero-shot). As this problem has not been fully explored in this field, we believe it can be further addressed in the near future.

### v3

@@ 0.B.2 Generalization Ability
We have shown the effectiveness of the continuous feature space when sufficient training data is available. However, in this subsection, we suppose a more challenging circumstance where the quantity of normal training data is significantly limited. When the normal images are not sufficient, models often suffer from learning the representative characteristics of the class. For instance, in discrete feature space, if an input feature significantly deviates from the normal features in the space, the model struggles to distinguish whether the distance stems from unseen normal or abnormal features since it has few normal features to be referenced. We compared CRAD with other methods by training each model with few-shot images during 100 epochs. As shown in Tab. 7, the methods with discrete feature space show limited performance under these constraints. On the other hand, CRAD allows continuous normal representation by interpolating the learned normal features in spite of the scarcity, resulting in superior performance under few-shot settings. This demonstrates the generalization ability of the continuous space even in a more challenging scenario. Despite outperforming the other baselines, CRAD’s performance under the extremely scarce dataset (i.e., 1-shot scenario) leaves room for improvement and we believe is not entirely satisfactory for practical uses. We recognize this as a limitation of CRAD, which will be discussed in Appendix 0.C.
@@
PatchCore UniAD CRAD PatchCore UniAD CRAD
1-shot 82.7 80.3 83.3 91.8 91.2 92.2
2-shot 87.5 82.2 88.2 93.8 92.4 95.1
4-shot 89.3 83.6 92.4 94.6 93.1 96.2
@@
Table 8: Ablation study on the grid dimension ($C_{l}$ and $C_{g}$) of local and global representations. The values in the parenthesis denote the grid resolution ($R_{l}$ and $R_{g}$) of each dimension. For instance, 2D (4) in the first row denotes that we set the grid dimension to 2 and each dimension’s resolution to 4. The performances are measured in AUROC on MVTec AD. CRAD’s setting is highlighted in bold.
@@ 0.B.3 More Ablation Studies
Grid dimension. We implemented CRAD with different grid dimensions (equivalent to the dimension of coordinates). As shown in Tab. 8, although the higher dimension of local and global representation slightly increases the detection and localization performance respectively, two representations both with 2-dimension are well balanced, resulting in the optimal performance. However, the results indicate CRAD is not sensitive to grid dimensions, which can be beneficial in a practical scenario.
Feature refinement. We conducted an ablation study on the feature refinement of CRAD, which employs both Mean Squared Error (MSE) and cosine similarity. As shown in Tab. 9, MSE and cosine similarity show comparable performance when they are applied individually. By combining the two metrics, CRAD achieves the best performance.
@@
0.B.4 Additional Dataset
We conducted an additional experiment using Real-IAD [realiad] (single view), a recently introduced and more challenging 30-class AD dataset. As shown in Tab. 10, CRAD outperforms UniAD by a large margin.
Method Det. (AUROC) Loc. (AUPRO)
UniAD 82.9 86.1
CRAD 88.6 87.5
Table 10: Performance evaluation using Real-IAD dataset in the unified setting.
0.B.5 More Qualitative Results
We provide more qualitative results of CRAD in Fig. 6.
Figure 6: Qualitative results of CRAD for each class of MVTec AD. Each row represents anomaly images, ground truths, prediction of UniAD, and prediction of CRAD.
@@ Appendix 0.C Limitation
Although CRAD demonstrates its superior generalization capabilities compared to existing methods, as shown in Sec. 0.B.2, it struggles when dealing with novel features. For instance, CRAD can generalize a sparse set of features in a class (i.e., few-shot) by interpolating within their feature space. However, this cannot be the case with extremely limited or no data (i.e., 1-shot or zero-shot). As this problem has not been fully explored in this field, we believe it can be further addressed in the near future.